# Roll in the Tanks! Measuring Left-wing Extremism on Reddit at Scale

### v3

@@ 5.1. What do tankies talk about?
Results. Table 2 presents the topics discussed by tankies ordered by the number of posts in each topic, as well as tankies’ RTPR for each topic. We find that the most popular topic among tankies is the Uyghur (a Turkic ethnic group from Xinjiang) Genocide occurring in Xinjiang, China, accounting for 1.11% of all tankies posts. In addition, we see that topics related to ideologies (socialism, communism, fascism, and capitalism) and two AES states (North Korea and USSR) in the top 10. When we look at the RTPRs, we find that they generally focus more on recent, state-level political events compared to other far-left communities. As we can see from the table, tankies have RTPR = 1 for the topics related to Uyghur genocide (#1), North Korea (#4), Russian invasion of Ukraine (#24), China-US conflicts (#29), Israel-Palestine conflict (#37), communism (#2), fascism (#8), Stalin (#17).
@@
Conversely, tankies rank in the bottom half when it comes to focusing on social problems. We find tankies have RTPRs between 4-7 on topics related to police (#20), climate change (#42), healthcare (#44), housing (#45), guns (#46), unions (#48), and taxes (#49). Additionally, our analysis reveals that tankies have RTPRs 2-3 on topics related to USSR (#9), tankies (#25), and Cuba (#31). However, even though topics related to socialism and capitalism are among the top 10 most popular topics of tankies, they have RTPRs of 6 and 5 for these topics, respectively. These results suggest that tankies prioritize discussions of state-level political events over the core ideological issues and social problems that have long concerned the far-left community.
5.2. What are their lexical differences?
So far, we discover that tankies focus more on state-level political events and less on social issues and ideological principles than other far-left communities. In this section, we investigate the differences in language use between tankies and other far-left communities. Given the diverse beliefs represented in left-wing extremism, conventional tools often fall short in effectively capturing their complexity (Jungkunz, 2019). To capture this nuance, we leverage the concept of alignment. This involves mapping tankies’ word usage to other far-left groups using multidimensional embeddings, providing a deeper understanding of their unique characteristics. We focus on two concepts:
For a more comprehensive overview of the methodology and preprocessing steps, please refer to Appendix E.
Discovering Misalignments & Conceptual Homomorphisms. We look for misaligned and conceptually homomorphic word pairs (with tankies as the source community) that have a cosine similarity greater than 0.5 (27% of all cosine similarites). This threshold has been previously adopted in studies using Word2Vec to achieve more reliable outcomes, demonstrating its effectiveness in various research contexts (Xue et al., 2021; Ge and Moh, 2017; Nguyen et al., 2020). The CDF of cosine similarities of the misaligned and conceptually homomorphic pairs between tankies and r/communism, r/socialism, r/DemocraticSocialism, and r/Anarchism can be seen in Figure 8 in Appendix E.
@@ 5.2. What are their lexical differences?
Views on US Politics (No Highlight). We find that tankies frequently discuss Joe Biden, in the same way that r/DemocraticSocialism do about Donald Trump. This tendency can be observed through the usage of phrases like “Bidens” & “Trumps” and “Sleepy Joe” & “Biden.” Additionally, the anti-democratic leaning of tankies is evident from their discussions on topics such as “popularity contest” & “elections.”
@@ 6. Discussion & Conclusion
Our analysis finds that the tankie community forms a distinct and cohesive group positioned at the periphery of the larger far-left spectrum on Reddit. We also observe that tankies gained increasing popularity within established far-left communities as their visibility and prominence grew. Further, we find notable differences in both the content and tone of tankie discussions compared to other far-left communities. Our topic analysis reveals that tankies prioritize state-level political events and discussions surrounding authoritarian regimes over core left-wing issues like healthcare and housing. Our lexical analysis shows consistent vocabulary misalignments and conceptual homomorphisms between tankies and other far-left communities, indicating tankies’ divergent ideological focus. Furthermore, we find tankies exhibit higher levels of toxicity, and they frequently share links from revisionist news sources.
@@ 7. Future Work
What’s in a name? As noted previously, there has definitely been a drift from the original meaning of “tankie” to the current meaning. One (perhaps cynical) explanation for this might be that online spaces have enabled a more fundamental shift in extremism. For example, it remains unclear if tankies’ online rhetoric translates to more traditionally explored aspects of leftist political movements like collective action, or if it is primarily a community engaged in extreme online contrarianism a la 4chan (Hine et al., 2017).
At the same time, the presence of the “tankie” label itself suggests that we have a relatively limited understanding of the contemporary far-left’s self-identification. While “tankie” is a succinct, relatively easy to understand label that has a relatively clear etymology, and Figure 5 shows that tankies form a clearly defined cluster of subreddits, the same figure also shows that the far-left is relatively diverse and there are many different ideologies with similar but meaningfully different labels (e.g., r/DemocraticSocialism and r/socialism).
@@ 8. Limitations
Tankies are almost completely unexplored in measurement-driven research, and relatively unexplored in the broader literature when compared to more “mainstream” far-left ideologies like communism. This raises some concerns with respect to our interpretations and our dataset itself. While our results are definitely in-line with what little literature exists on the topic, the name “tankies” has absolutely drifted from its original meaning. We rely in large part on the tankie community on Reddit’s self description of “tankie,” and although it aligns with recent use in the literature (Petterson, 2020), it is worth recognizing that there is not a rich body of theoretical work for us to draw from when interpreting results. Our study focuses on six tankie subreddits, with a substantial portion of the data coming from r/GenZedong. This focus may inherently emphasize the perspectives and behaviors of a specific subset within the broader tankie community. Additionally, smaller far-left subreddits are not included in our comparison with tankies. While these subreddits could offer alternative perspectives, we believe that the prominence and activity levels of the larger subreddits provide a strong and reliable foundation for studying the tankie community on Reddit.
@@
9. Acknowledgments
This material is based upon work supported by the European Union under the project MedDMO (Grant Agreement no. 101083756).
@@ B.1. Tankie Subreddits
r/GenZhou is the fourth most popular tankie subreddit, with 28,606 posts from 3,620 authors since December 2020. r/GenZhou refers not only to leftist subreddits, but also subreddits that are ideologically opposed to tankies (r/tankiejerk, r/askaconservative, r/centrist, and r/nato) as well as debate subreddits (r/answers and r/PoliticalDiscussion). One explanation might be that r/GenZhou could be pointing these communities to their user-base to call for support in certain discussions. We also find r/GenZhou referring to r/AskARussian, which might reflect a pro-Russian leaning. Reddit banned r/GenZhou on April 4, 2022 (Reddit, 2022) for “creating or repurposing a sub to reconstitute or serve the same objective as a previously banned or quarantined subreddit” (i.e., ban evasion).
@@ B.2. Network Metrics
r/communism r/socialism r/DemocraticSocialism r/Anarchism Named Entity #Posts % Named Entity #Posts % Named Entity #Posts % Named Entity #Posts % China 7,157 4.15 US 17,172 3.33 Biden 13,821 3.45 One 13,263 2.56 US 5,635 3.27 One 12,362 2.40 Trump 11,000 2.74 First 11,879 2.29 One 4,162 2.41 First 11,356 2.20 Bernie 10,281 2.56 US 10,038 1.93 First 3,959 2.29 China 10,491 2.03 US 8,616 2.15 Two 6,737 1.30 USSR 2,942 1.70 American 7,573 1.47 One 8,564 2.13 American 5,012 0.96 Marxist 2,895 1.68 Two 6,740 1.30 First 7,200 1.79 China 4,770 0.92 Chinese 2,649 1.53 America 6,578 1.27 democrats 7,116 1.77 Today 4,298 0.83 American 2,597 1.50 USSR 5,478 1.06 Two 5,916 1.47 America 4,164 0.80 Marx 2,333 1.35 Marx 5,470 1.06 America 5,881 1.46 2 3,325 0.64 Two 2,331 1.35 Today 5,096 0.98 Republicans 5,758 1.43 Second 2,937 0.56 Stalin 2,295 1.33 Russia 4,675 0.90 American 5,316 1.32 1 2,864 0.55 Lenin 2,238 1.29 Cuba 4,457 0.86 Americans 3,720 0.92 Reddit 2,858 0.55 Today 1,956 1.13 Marxist 4,283 0.83 Republican 3,365 0.84 Trump 2,846 0.54 Cuba 1,837 1.06 Socialists 4,072 0.79 DNC 3,009 0.75 Russia 2,758 0.53 America 1,761 1.02 Lenin 4,065 0.78 2 2,928 0.73 Marx 2,519 0.48 Russia 1,653 0.95 Stalin 3,701 0.71 Democratic 2,922 0.72 Nazis 2,469 0.47 Mao 1,592 0.92 Trump 3,666 0.71 Democrat 2,760 0.68 Anarchists 2,298 0.44 Soviet 1,532 0.88 2 3,600 0.69 Today 2,481 0.61 3 2,224 0.42 Soviet Union 1,230 0.71 Chinese 3,322 0.64 1 2,331 0.58 USSR 2,007 0.38 Communist 1,224 0.71 Second 3,258 0.63 Obama 2,316 0.57 Biden 1,947 0.37 r/alltheleft r/Marxism r/IWW Named Entity #Posts % Named Entity #Posts % Named Entity #Posts % US 2,328 1.91 Marx 3,297 12.91 IWW 2,932 10.2 Biden 2,188 1.79 Marxist 1,817 7.11 One 903 3.16 One 2,016 1.65 One 1,203 4.71 First 678 2.37 Trump 1,947 1.60 First 1,169 4.57 US 439 1.53 First 1,744 1.43 Two 754 2.95 Two 387 1.35 Democrats 1,430 1.17 Marxists 730 2.85 Solidarity 354 1.23 American 1,394 1.14 Lenin 670 2.62 Today 273 0.95 America 1,272 1.04 US 588 2.30 1 219 0.76 Two 1,187 0.97 China 566 2.21 2 217 0.75 100% 1,168 0.95 Today 547 2.14 American 193 0.67 China 994 0.81 USSR 475 1.86 GEB 161 0.56 Republicans 942 0.77 Engels 474 1.85 Second 159 0.55 Second 780 0.64 Stalin 452 1.77 3 155 0.54 Bernie 748 0.61 1 447 1.75 Wobblies 151 0.52 Americans 680 0.55 2 389 1.52 Amazon 145 0.50 Obama 623 0.51 Second 337 1.32 DSA 131 0.45 Today 621 0.51 3 284 1.11 Reddit 130 0.45 shitpisscum1312 595 0.48 American 282 1.10 America 128 0.44 Fucking cooldownbot 595 0.48 Russia 223 0.87 NLRB 124 0.43 Russia 572 0.47 German 219 0.85 Biden 121 0.42
Table 8. Top 20 named entities of far-left communities other than tankies.
@@ B.3. User Migrations Over Time
During the first month after r/GenZedong was founded, 58.4% of the tankies user-base came from other far-left communities. On average, 34.3% of tankies MAU came from other far-left communities during the first period. However, during the second period, there was a clear decrease in migrations from other far-left communities to tankies, with a mean MRD of 17.3%. The highest migrations to tankies come from r/communism and r/socialism for most of the months in our dataset. During the first period, the mean MRDs from these communities to tankies are 15.8%. The second period also saw a decrease in MRDs from these communities, with mean MRDs of 5.7% and 8.4% respectively.
We find strong negative correlations between the monthly user counts of tankies and MRDs to tankies from r/DemocraticSocialism ($\rho=-0.71$) and r/Anarchism ($\rho=-0.71$). We also find moderate negative correlations for r/Marxism ($\rho=-0.69$), r/communism ($\rho=-0.68$), r/socialism ($\rho=-0.64$), and r/IWW ($\rho=-0.60$). In contrast, we find a moderate positive correlation for r/alltheleft ($\rho=0.49$). We do not observe any trend for this community’s MRDs to tankies, and we find decreasing trends for MRDs from other far-left communities to tankies when applying the Mann-Kendall test.

### v4

@@ 5.1. What do tankies talk about?
Figure 7. Monthly proportion of posts in highlighted topics assigned to state-level politics and core social issues for tankies and other far-left communities. Tankies consistently devote a larger share of these posts to state-level politics, while other far-left communities focus more on social issues.
@@
Results. Table 2 presents the topics discussed by tankies ordered by the number of posts in each topic, as well as tankies’ RTPR for each topic. We find that the most popular topic among tankies is the Uyghur (a Turkic ethnic group from Xinjiang) Genocide occurring in Xinjiang, China, accounting for 1.11% of all tankies posts. In addition, we see that topics related to ideologies (socialism, communism, fascism, and capitalism) and two AES states (North Korea and USSR) in the top 10. When we look at the RTPRs, we find that they generally focus more on recent, state-level political events compared to other far-left communities (see highlighted topics in Table 2). As we can see from the table, tankies have RTPR = 1 for the topics related to Uyghur genocide (#1), North Korea (#4), Russian invasion of Ukraine (#24), China-US conflicts (#29), Israel-Palestine conflict (#37), communism (#2), fascism (#8), Stalin (#17).
Conversely, tankies rank in the bottom half when it comes to focusing on social problems. We find tankies have RTPRs between 4-7 on topics related to police (#20), climate change (#42), healthcare (#44), housing (#45), guns (#46), unions (#48), and taxes (#49). Additionally, our analysis reveals that tankies have RTPRs 2-3 on topics related to USSR (#9), tankies (#25), and Cuba (#31).
Figure 7 plots the monthly proportion of posts in the highlighted topics for tankies and other far-left communities. We find that tankies consistently prioritize state-level politics, with these topics typically accounting for 60-70% of the highlighted topic posts, and social issues making up the remaining 30-40%. In contrast, other far-left communities show the opposite pattern, where between 70-80% of the highlighted topic posts concern core social issues, while only about 20-30% focus on state-level politics. Notably, we observe a visible increase in attention to state-level politcs in February-March 2022, coinciding with the Russian invasion of Ukraine, but the relative ordering between tankies and other far-left communities remains unchanged.
These results suggest that tankies prioritize discussions of state-level political events over the core ideological issues and social problems that have long concerned the far-left community.
5.2. What are their lexical differences?
So far, we discover that tankies focus more on state-level political events and less on social issues and ideological principles than other far-left communities. In this section, we investigate the differences in language use between tankies and other far-left communities. Given the diverse beliefs represented in left-wing extremism, conventional tools often fall short in effectively capturing their complexity (Jungkunz, 2019). To capture this nuance, we leverage the concept of alignment. This involves mapping tankies’ word usage to other far-left groups using multidimensional embeddings, providing a deeper understanding of their unique characteristics.
Aligning Multidimensional Word Embeddings. Understanding the language usage in polarized communities is challenging due to their distinct vocabularies for identical concepts. E.g., Democrats and Republicans are political adversaries. Thus, a Democrat might say that “Republicans are America’s greatest threat” (and vice versa), making direct comparisons of their language less useful when it comes to polarizing topics. To address this, KhudaBukhsh et al. (KhudaBukhsh et al., 2021) formulate the problem as a form of translation between two groups using the same natural language. While not every “translation” is as clear-cut as the previous toy example, this technique has seen refinement (Milbauer et al., 2021) to capture more complex, multidimensional aspects of ideological differences. As such, it emerges as an ideal tool for distinguishing disparities between tankies and other left-wing ideologies.
In a nutshell, we align two multidimensional word embeddings by creating a bilingual lexicon to anchor the alignment. We then use a translation function to project the embeddings from one community to another, and once both vocabularies live in the same vector space, we compute cosine similarity between corresponding word vectors to quantify how similarly each term is used across communities. We focus on two concepts:
Discovering Misalignments & Conceptual Homomorphisms. We look for misaligned and conceptually homomorphic word pairs (with tankies as the source community) that have a cosine similarity greater than 0.5 (27% of all cosine similarities). This threshold has been previously adopted in studies using Word2Vec to achieve more reliable outcomes, demonstrating its effectiveness in various research contexts (Xue et al., 2021; Ge and Moh, 2017; Nguyen et al., 2020). The CDF of cosine similarities of the misaligned and conceptually homomorphic pairs between tankies and r/communism, r/socialism, r/DemocraticSocialism, and r/Anarchism can be seen in Figure 10 in Appendix E.
@@ 5.2. What are their lexical differences?
Views on US Politics (No Highlight). We find that tankies frequently discuss Joe Biden, in the same way that r/DemocraticSocialism do about Donald Trump. This tendency can be observed through the usage of phrases like “Bidens” & “Trumps” and “Sleepy Joe” & “Biden.” Additionally, the anti-democratic leaning of tankies is evident from their discussions on topics, e.g., “popularity contest” & “elections.”
@@ 6. Discussion & Conclusion
Our analysis finds that the tankie community forms a distinct and cohesive group positioned at the periphery of the larger far-left spectrum on Reddit. We also observe that tankies gained increasing popularity within established far-left communities as their visibility and prominence grew. Further, notable differences in both the content and tone of tankie discussions compared to other far-left communities. Our topic analysis reveals that tankies prioritize state-level political events and discussions surrounding authoritarian regimes over core left-wing issues like healthcare and housing. Our lexical analysis shows consistent vocabulary misalignments and conceptual homomorphisms between tankies and other far-left communities, indicating tankies’ divergent ideological focus. Furthermore, we find tankies exhibit higher levels of toxicity, and they frequently share links from revisionist news sources.
@@ 7. Future Work
Sentiment and topic dynamics. Future work could more explicitly link sentiment measures to specific topics, highlighting how different communities not only discuss the same issues but also diverge in their attitudes toward them. In parallel, examining tankies’ engagement with non-political, everyday topics would help situate their extremist rhetoric within the broader texture of their online identity.
What’s in a name? As noted previously, there has definitely been a drift from the original meaning of “tankie” to the current meaning. One (perhaps cynical) explanation for this might be that online spaces have enabled a more fundamental shift in extremism. For example, it remains unclear if tankies’ online rhetoric translates to more traditionally explored aspects of leftist political movements like collective action, or if it is primarily a community engaged in extreme online contrarianism à la 4chan (Hine et al., 2017).
At the same time, the presence of the “tankie” label itself suggests that we have a relatively limited understanding of the contemporary far-left’s self-identification While “tankie” is a succinct, relatively easy to understand label that has a relatively clear etymology, and Figure 6(a) shows that tankies form a clearly defined cluster of subreddits, the same figure also shows that the far-left is relatively diverse and there are many different ideologies with similar but meaningfully different labels (e.g., r/DemocraticSocialism and r/SocialDemocracy).
@@ 8. Limitations
Tankies are almost completely unexplored in measurement-driven research, and relatively unexplored in the broader literature when compared to more “mainstream” far-left ideologies like communism. This raises some concerns with respect to our interpretations and our dataset itself. While our results are definitely in-line with what little literature exists on the topic, the name “tankies” has absolutely drifted from its original meaning. We rely in large part on the tankie community on Reddit’s self description of “tankie,” and although it aligns with recent use in the literature (Petterson, 2020), it is worth recognizing that there is not a rich body of theoretical work for us to draw from when interpreting results. Our study focuses on six tankie subreddits, with a substantial portion of the data coming from r/GenZedong. This focus may inherently emphasize the perspectives and behaviors of a specific subset within the broader tankie community. Additionally, smaller far-left subreddits are not included in our comparison with tankies. While these subreddits could offer alternative perspectives, we believe that the prominence and activity levels of the larger subreddits provide a strong and reliable foundation for studying the tankie community on Reddit. Additionally, our analysis is restricted to 2019-2022, which corresponds to the lifespan of the largest self-identified tankie subreddit on Reddit. This means we do not capture earlier phases of tankies on other platforms or the evolution of successor communities after 2022, which we leave for future work.
@@ B.1. Tankie Subreddits
r/GenZhou is the fourth most popular tankie subreddit, with 28,606 posts from 3,620 authors since December 2020. r/GenZhou refers not only to leftist subreddits, but also subreddits that are ideologically opposed to tankies (r/tankiejerk, r/askaconservative, r/centrist, and r/nato) as well as debate subreddits (r/answers and r/PoliticalDiscussion). One explanation might be that r/GenZhou could be pointing these communities to their user base to call for support in certain discussions. We also find r/GenZhou referring to r/AskARussian, which might reflect a pro-Russian leaning. Reddit banned r/GenZhou on April 4, 2022 (Reddit, 2022) for “creating or repurposing a sub to reconstitute or serve the same objective as a previously banned or quarantined subreddit” (i.e., ban evasion).
@@ B.2. Network Metrics
r/communism r/socialism r/DemocraticSocialism r/Anarchism Named Entity #Posts % Named Entity #Posts % Named Entity #Posts % Named Entity #Posts % China 7,157 4.15 US 17,172 3.33 Biden 13,821 3.45 One 13,263 2.56 US 5,635 3.27 One 12,362 2.40 Trump 11,000 2.74 First 11,879 2.29 One 4,162 2.41 First 11,356 2.20 Bernie 10,281 2.56 US 10,038 1.93 First 3,959 2.29 China 10,491 2.03 US 8,616 2.15 Two 6,737 1.30 USSR 2,942 1.70 American 7,573 1.47 One 8,564 2.13 American 5,012 0.96 Marxist 2,895 1.68 Two 6,740 1.30 First 7,200 1.79 China 4,770 0.92 Chinese 2,649 1.53 America 6,578 1.27 democrats 7,116 1.77 Today 4,298 0.83 American 2,597 1.50 USSR 5,478 1.06 Two 5,916 1.47 America 4,164 0.80 Marx 2,333 1.35 Marx 5,470 1.06 America 5,881 1.46 2 3,325 0.64 Two 2,331 1.35 Today 5,096 0.98 Republicans 5,758 1.43 Second 2,937 0.56 Two 2,331 1.35 Today 5,096 0.98 Republicans 5,758 1.43 Second 2,937 0.56 Stalin 2,295 1.33 Russia 4,675 0.90 American 5,316 1.32 1 2,864 0.55 Lenin 2,238 1.29 Cuba 4,457 0.86 Americans 3,720 0.92 Reddit 2,858 0.55 Today 1,956 1.13 Marxist 4,283 0.83 Republican 3,365 0.84 Trump 2,846 0.54 Cuba 1,837 1.06 Socialists 4,072 0.79 DNC 3,009 0.75 Russia 2,758 0.53 America 1,761 1.02 Lenin 4,065 0.78 2 2,928 0.73 Marx 2,519 0.48 Russia 1,653 0.95 Stalin 3,701 0.71 Democratic 2,922 0.72 Nazis 2,469 0.47 Mao 1,592 0.92 Trump 3,666 0.71 Democrat 2,760 0.68 Anarchists 2,298 0.44 Soviet 1,532 0.88 2 3,600 0.69 Today 2,481 0.61 3 2,224 0.42 Soviet Union 1,230 0.71 Chinese 3,322 0.64 1 2,331 0.58 USSR 2,007 0.38 Communist 1,224 0.71 Second 3,258 0.63 Obama 2,316 0.57 Biden 1,947 0.37 r/alltheleft r/Marxism r/IWW Named Entity #Posts % Named Entity #Posts % Named Entity #Posts % US 2,328 1.91 Marx 3,297 12.91 IWW 2,932 10.2 Biden 2,188 1.79 Marxist 1,817 7.11 One 903 3.16 One 2,016 1.65 One 1,203 4.71 First 678 2.37 Trump 1,947 1.60 First 1,169 4.57 US 439 1.53 First 1,744 1.43 Two 754 2.95 Two 387 1.35 Democrats 1,430 1.17 Marxists 730 2.85 Solidarity 354 1.23 American 1,394 1.14 Lenin 670 2.62 Today 273 0.95 America 1,272 1.04 US 588 2.30 1 219 0.76 Two 1,187 0.97 China 566 2.21 2 217 0.75 100% 1,168 0.95 Today 547 2.14 American 193 0.67 China 994 0.81 USSR 475 1.86 GEB 161 0.56 Republicans 942 0.77 Engels 474 1.85 Second 159 0.55 Second 780 0.64 Stalin 452 1.77 3 155 0.54 Bernie 748 0.61 1 447 1.75 Wobblies 151 0.52 Americans 680 0.55 2 389 1.52 Amazon 145 0.50 Obama 623 0.51 Second 337 1.32 DSA 131 0.45 Today 621 0.51 3 284 1.11 Reddit 130 0.45 shitpisscum1312 595 0.48 American 282 1.10 America 128 0.44 Fucking cooldownbot 595 0.48 Russia 223 0.87 NLRB 124 0.43 Russia 572 0.47 German 219 0.85 Biden 121 0.42
Table 8. Top 20 named entities of far-left communities other than tankies
@@ B.3. User Migrations Over Time
During the second month after r/GenZedong was founded, 32.4% of the tankies user base came from other far-left communities. On average, 27.7% of tankies MAU came from other far-left communities during the first period. However, during the second period, there was a clear decrease in migrations from other far-left communities to tankies, with a mean MRD of 14.9%. The highest migrations to tankies come from r/communism and r/socialism for most of the months in our dataset. During the first period, the mean MRDs from these communities to tankies are 13%. The second period also saw a decrease in MRDs from these communities, with mean MRDs of 4.8% and 7.4% respectively.
We find strong negative correlations between the monthly user counts of tankies and MRDs to tankies from r/DemocraticSocialism ($\rho=-0.78$), r/Anarchism ($\rho=-0.78$), r/communism ($\rho=-0.74$), r/socialism ($\rho=-0.72$), and r/Marxism ($\rho=-0.72$). We also find a moderate negative correlation for r/IWW ($\rho=-0.67$). In contrast, we find a moderate positive correlation for r/alltheleft ($\rho=0.48$). We do not observe any trend for this community’s MRDs to tankies, and we find decreasing trends for MRDs from other far-left communities to tankies when applying the Mann-Kendall test.